OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 16 times y cubed − the square root of y cubed 3 times y times the square root of y 3 times y squared times the square root of y 4 times y squared times the square root of y 4 times y times the square root of y

OpenStudy (anonymous):

\[\sqrt{16y^3} - \sqrt{y^3}\]

OpenStudy (agent0smith):

\[\Large \sqrt { 16 y^3} - \sqrt {y^3}\] Break up the square roots like last time \[\Large \sqrt { 16} \sqrt{y^3} - \sqrt {y^3}\]

OpenStudy (anonymous):

4 sqrt y^3-y^3

OpenStudy (anonymous):

now what do you do?

OpenStudy (agent0smith):

Now break the y^3 up into y^2*y... \[\large y^3 = y^2 \times y\] understand why? \[\Large \sqrt { 16} \sqrt{y^3} - \sqrt {y^3} = \] \[\Large 4 \sqrt{y^2 \times y} - \sqrt {y^2 \times y} = \]

OpenStudy (anonymous):

yes, and when you do that you get \[4y \sqrt{y}\] ?

OpenStudy (agent0smith):

I think you missed a term, check again - you're partially correct :)

OpenStudy (anonymous):

oh I got it now! \[4y^2\sqrt{y}\]

OpenStudy (agent0smith):

Hmm, not quite. Do it one step at a time, simplify the left term and the right term \[\Large 4 \sqrt{y^2 \times y} - \sqrt {y^2 \times y} =\] \[\Large 4 y \sqrt{ y} - y\sqrt { y} =\]

OpenStudy (anonymous):

3y sqrt y

OpenStudy (agent0smith):

Excellent :)

OpenStudy (anonymous):

haha, thank you :)

OpenStudy (agent0smith):

Welcome :)